OpenStudy (anonymous):

PLEASE SOLVE THE FOLLOWING SYSTEM OF EQUATIONS. 3.2x-0.6y=8 -0.8x+0.9y=4

OpenStudy (anonymous):

is the first one x=0.1875y+2.5 ?

OpenStudy (cwrw238):

you can solve these by elimination method multiply the second equation by 4 to get the 2 equations 3.2x -0.6y = 8 -3.2x + 3.6y = 16

OpenStudy (anonymous):

am i suppose to solve both at the same time?

OpenStudy (cwrw238):

now if you add these 2 equations you eliminate x then solve for y having found the value of y substitute it in one of the original equations to find value of x

OpenStudy (cwrw238):

do you follow that?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

do i have to solve each separately?

OpenStudy (anonymous):

was my first answer correct?

OpenStudy (cwrw238):

no you solve them the way I've just explained 3.2x -0.6y = 8 -3.2x + 3.6y = 16 add these 2 equations ( note -3.2x + 3.2x = 0) 0 + 3y = 24 3y = 24 y = 24/3 = 8 now plug in y = 8 into the first equation and solve for x

OpenStudy (cwrw238):

the answer you gave first was the value of x in terms of y the question asks you to solve the system of equations that is find the value of x and y I've done half of it for you ( y = 80 now you need to find the value of x as I explained in my last post.

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

y=80 is the answer for one?

OpenStudy (anonymous):